# but Conditions Apply

### v1

@@ 1 Introduction
Models like Stable Diffusion Esser et al. (2024); Rombach et al. (2022) have been trained on billions of image-text pairs and can generate highly detailed and coherent images that match textual descriptions. Their ever-increasing ability to generate complex compositional scenes Huang et al. (2023); Ghosh et al. (2023) suggests they have developed a strong understanding of image-text relationships and can effectively align visual and textual concepts. Diffusion models are trained with pixel-wise supervision, so they may be less prone to learn shortcuts, compared to discriminatively-trained models like CLIP Radford et al. (2021), which often are insensitive to word order Yuksekgonul et al. (2023); Hsieh et al. (2023); Dumpala et al. (2024), struggle with spatial relationships, counting Tong et al. (2024); Paiss et al. (2023), and compositions Thrush et al. (2022); Koishigarina et al. (2025); Uselis et al. . It is thus natural to ask: can we transfer the compositional capabilities of generative models to discriminative compositional tasks?
There is growing interest in leveraging the strong generative capabilities of diffusion models for broader discriminative tasks such as classification, shape-texture bias, and depth estimation Zhao et al. (2023); Saxena et al. (2023); Jaini et al. (2023). One emerging approach, known as Diffusion Classifiers, repurposes diffusion models for zero-shot classification using their original training objective Li et al. (2023); Krojer et al. (2023); Clark and Jaini (2023); He et al. (2024); Jaini et al. (2023). Notably, they have outperformed CLIP on compositional benchmarks like CLEVR Johnson et al. (2017) and Winoground Thrush et al. (2022), which require reasoning over multiple objects and attributes. However, existing studies are limited in scope, often relying on a small number of benchmarks and lacking a systematic analysis.
@@
Hypothesis 3: The domain gap can be mitigated by timestep weighting. Lastly, we investigate how diffusion timesteps impact discriminative performance by optimizing timestep weights for downstream tasks. Prior work has explored how diffusion models generate images through a structured progression across timesteps Li et al. ; Wang and Vastola (2023). However, in diffusion classifiers, either uniform timesteps or fixed timestep weightings are typically used across all datasets and models—an area that remains underexplored. In contrast, we find that SD3-m is particularly sensitive to timestep selection. Our results show that even low-shot timestep tuning (using just 5% of the target dataset can significantly mitigate the performance drop of SD3-m on real-world compositional benchmarks. We hypothesize that SD3-m’s timestep sensitivity is closely linked to its susceptibility to the domain gap. To further examine this relationship, we incorporate CLIP-based image encoders to quantify visual similarity between domains, and analyze how it correlates with optimal timestep weighting. We find that timestep weighting is especially effective in scenarios with large domain gaps.
@@ 2 Related work
Diffusion models. Generative models have demonstrated impressive performance in producing realistic images Ramesh et al. (2022); Yang et al. (2024); Chen et al. (2023), videos Blattmann et al. (2023); Zheng et al. (2024), and audio Evans et al. (2024); Liu et al. (2023a); Jeong et al. (2025). In particular, text-to-image diffusion models Ho et al. (2020) iteratively refine images conditioning on text prompts by adding and removing noise, achieving remarkable quality. A widely used open-source example is Stable Diffusion (SD) Rombach et al. (2022). Earlier SD versions Rombach et al. (2022) are based on a UNet Ronneberger et al. (2015) backbone, incorporating ResNet blocks He et al. (2016) and attention mechanisms Vaswani et al. (2017). With the rise of transformer-based designs Vaswani et al. (2017), the latest Stable Diffusion 3 series Esser et al. (2024) adopts this architecture, further boosting performance. It also introduces new noise sampling strategies for training Rectified Flow models Liu et al. (2023b); Albergo and Vanden-Eijnden (2023); Lipman et al. (2023). Our analysis focuses on Stable Diffusion versions 1.5, 2.0, and 3-m, examining their architectural evolution and performance across scales.
@@ 3.1 Preliminaries: diffusion classifiers
where $w_{t}$ are timestep weights, $\epsilon_{\Theta}$ is a neural network that predicts the noise $\boldsymbol{\epsilon}$ added to the latent $\mathbf{z}$ at timestep $t\in T$, and $\mathbf{c}$ is a conditioning text prompt. This loss is related to the ELBO of the conditional likelihood $p(\mathbf{z}|y)$, where $y$ is the class label, which allows us to use diffusion models for classification, as shown in Li et al. (2023); Krojer et al. (2023); Clark and Jaini (2023):
@@ 3.2 SD3-m as a classifier
SD3-m is a rectified flow model Liu et al. (2023b); Albergo and Vanden-Eijnden (2023) trained with a conditional flow matching (CFM) loss Lipman et al. (2023), which differs from the standard diffusion objectives used in SD1.5 and SD2.0. As a result, we cannot directly apply the same classifier objective used in earlier versions. By reparameterizing the CFM objective as a noise-prediction loss (see, e.g., Esser et al. (2024)), we can obtain
@@
Using this formulation, we can use SD3 as classifiers, despite its different underlying architecture. The only difference lies in the weighting function $w_{t}$, which for SD3 follows a logit-normal distribution rather than the uniform weighting used in SD1.5/2.0. However, we empirically find that uniform weighting performs better for classification. Details are given in Section A.1.
@@ 4 Self-Bench: a diagnostic benchmark
As shown in Figure 2 (left), existing benchmarks are diverse in terms of domains. However, we empirically observe that the generation results of diffusion models are not as diverse, having a preferred “native” style. For example, in SD3-m, objects are usually well-centered/focused and have a glossy aesthetic touch; images are in high resolution (see Figure 2, right). Additional examples in Figure A.7 (Supplemental) further support that SD3-m consistently produces images with a similar style.
@@
Figure 3: Diagnosing with Self-Bench. (i) Using Geneval’s prompts from six types, generate images. (ii) For each generated image, create discriminative tasks within its type from the prompts used in the generation process. (iii) Given the generated images (filtered by humans) and the discriminative tasks, benchmark the diffusion classifier.
@@ 5.1 Experimental setting
Stable Diffusion baselines. For the evaluation of diffusion classifiers, we use three versions of Stable Diffusion models: SD1.5, SD2.0 Rombach et al. (2022), and SD3-m Esser et al. (2024); each new version increases the number of parameters. We selected the baselines based on specific criteria, which are detailed in Section A.5.1 of the Supplemental. Although we also considered distilled variants of diffusion models (e.g., SDXL-Turbo Sauer et al. (2024)), we excluded them from our main evaluation due to their architectural and generative differences. Additional analysis is provided in Section A.7 of the Supplemental. For SD1.5 and SD2.0, we use the Euler Discrete scheduler Karras et al. (2022) and uniformly sample the timesteps. For SD3-m, we use the Flow Matching EulerDiscrete scheduler Esser et al. (2024) for flow matching diffusions, which was designed specifically for the SD3 series. We sample 30 timesteps from each model uniformly, following He et al. (2024). (The effect of different numbers of timesteps is discussed in Table A.2 in the Supplemental.)
CLIP baselines. We use vanilla CLIP models as key baselines for comparison on discriminative tasks. We use five different versions of CLIP models: RN50x64, ViT-B/32, ViT/L14, ViT/H14, and ViT/g14. We follow OpenAI’s implementation for the first three: RN50x64, ViT/B32, and ViT/L14 Radford et al. (2021). On the other hand, we follow OpenCLIP’s implementation for ViT-H/14 and ViT-g/14 Ilharco et al. (2021).
Metrics. The benchmarks vary in structure. Some present paired image-text inputs (i.e., two images and two prompts), while others use a single image with multiple candidate prompts. Across all setups, we primarily evaluate using image-to-text retrieval accuracy, which measures whether the model assigns the highest score to the correct prompt given an image. For paired settings, we compute retrieval accuracy based on whether the matching image-prompt pairs are correctly ranked relative to distractors. Further details on task-specific evaluation protocols are provided in Section A.5.2.
@@ 5.2 Scaling evaluation to ten benchmarks
Benchmarks. Our analysis incorporates ten benchmarks: Vismin Awal et al. (2024), EQBench Wang et al. (2023), MMVP Tong et al. (2024), CLEVR Johnson et al. (2017), Whatsup Kamath et al. (2023), Spec Peng et al. (2024), ARO Yuksekgonul et al. (2023), Sugarcrepe Hsieh et al. (2023), COLA Ray et al. (2023), and Winoground Thrush et al. (2022). Each benchmark contains different tasks. For example, Vismin includes tasks related to Object, Attribute, Position, and Counting, whereas COLA is focused on Attribute tasks. The left block in Figure 2 presents examples of these benchmarks. Further details on how the tasks are classified within our compositional task categories are provided in Table A.3 in the Supplemental. In total, we analyze 33 tasks in our main study.
@@
Results. Figure 4 shows the average performance of diffusion classifiers on the compositional benchmarks (see Figure A.10 in the Supplemental for the complete results). [^2] Tables A.10A.11A.12A.13A.14A.15A.16A.17A.18 and A.19 report the quantitative results for all benchmarks used in our evaluation. For the Position task, SD3-m performs the best compared to other diffusion models and CLIP models. However, in other tasks, CLIP models usually show better performance than diffusion classifiers, contrary to the findings of previous works Li et al. (2023); Clark and Jaini (2023). Surprisingly, among diffusion classifiers, SD3-m is not the best; often SD1.5 or SD2.0 models show better results. These results motivate us to deepen our analysis of when and why diffusion classifiers may underperform.
@@ 5.3 Studying domain effects via Self-Bench
In-domain evaluation. First, we examine how diffusion classifiers behave in-domain, which may serve as an upper-bound performance estimate, as domain effects are isolated. As shown in Figure 5, they perform better on Correct (yellow bars) than on Full (blue bars) samples. This indicates that the classifiers do not simply choose prompts used for generation; rather, they indeed distinguish between the correct and incorrect prompts.
@@
Figure 6 illustrates the average accuracy drop when moving from in-domain to cross-domain evaluation on Self-Bench. We observe accuracy degradation across all tasks, with SD3-m showing the most drop. While part of the performance gap may be due to task-specific weaknesses in each model, we assume the domain gaps, such as different data distribution, also play a key role in this degradation. Full results can be found in Figure A.6 in the Appendix.
@@ 5.4 Timestep weighting effects
Previous works have shown that diffusion models generate images from coarse to fine details over timesteps Li et al. ; Wang and Vastola (2023). However, in classification settings, it is still unclear how different noise levels (i.e., timesteps) affect performance across tasks (e.g., object recognition vs. attribute binding) or different domains (e.g., image style). Diffusion classifiers typically use uniform timestep weighting Li et al. (2023); Krojer et al. (2023) or a fixed timestep weighting scheme (e.g., $w_{t}=\exp(-7t)$Clark and Jaini (2023); Jaini et al. (2023) across all models (e.g., Imagen Saharia et al. (2022) and SD). We hypothesize that neither strategy is universally optimal (see ablation studies of uniform weighting and fixed timestep weighting in Sec. A.9 in the Supplemental) and that timestep weighting should be adapted to the model and task differently. Here, we investigate how different timesteps contribute to classification decisions. (Figure 10 in the Supplemental provides an intuitive explanation of why timesteps matter from a generative perspective.)
@@
To study this further, we conduct an experiment to approximate the effect of domain differences. We generate images using the original prompts from real-world compositional benchmarks. This yields two image sets for each task: (i) the original real-world dataset, and (ii) a synthetic variant generated using the same prompts. Both sets target the same task but differ in visual domain. Using a CLIP image encoder (ViT-B/32) Radford et al. (2021), we aim to capture the domain gap [^3] Indeed CLIP may capture both stylistic as well as semantic shifts, broadly referred to as ”domain.” between the two, by computing the L2 distance between average embeddings with randomly sampled 50 images. As shown in Figure 9, larger CLIP embedding distances (i.e., greater domain gap) are associated with greater reliance on timestep weighting, but only for SD3. This suggests a positive correlation between domain gap and the effectiveness of timestep weighting for SD3. In contrast, SD1 and SD2 do not exhibit such a trend. We hypothesize that, as shown in Figure 7, most timesteps in SD2 (also in SD1) are already effective, resulting in limited benefit from reweighting and thus obscuring any correlation with the domain gap. Details are in Section A.12 of the Appendix.
@@
To build intuition for how timestep selection impacts classification, we visualise generations from the SD3-m model in Figure 10, starting from a Self-Bench image generated by SD2.0 (“a parking meter and a teddy bear”). For each timestep $t$, we corrupt the original image with Gaussian noise corresponding to $t$, and then generate an image by running the diffusion model for 20 denoising steps from $t$ down to 00, conditioning on either a correct or incorrect prompt. We set the classifier-free guidance coefficient to 0.0 (i.e., using only the conditional prompt) to match the discriminative setting used in classification with diffusion.

### v3

@@ 1 Introduction
Models like Stable Diffusion Esser et al. (2024); Rombach et al. (2022) have been trained on billions of image-text pairs and can generate highly detailed and coherent images that match textual descriptions. Their ever-increasing ability to generate complex compositional scenes Huang et al. (2023); Ghosh et al. (2023) suggests they have developed a strong understanding of image-text relationships and can effectively align visual and textual concepts. Diffusion models are trained with pixel-wise supervision, so they may be less prone to learn shortcuts, compared to discriminatively-trained models like CLIP Radford et al. (2021), which often are insensitive to word order Yuksekgonul et al. (2023); Hsieh et al. (2023); Dumpala et al. (2024), struggle with spatial relationships, counting Tong et al. (2024); Paiss et al. (2023), and compositions Thrush et al. (2022); Koishigarina et al. (2025); Uselis et al. ; Uselis et al. (2025). It is thus natural to ask: can we transfer the compositional capabilities of generative models to discriminative compositional tasks?
There is growing interest in leveraging the strong generative capabilities of diffusion models for broader discriminative tasks such as classification, shape-texture bias, and depth estimation. Two main approaches have emerged: one line of work treats diffusion models as feature extractors, training task-specific classifiers in a supervised manner Zhao et al. (2023); Saxena et al. (2023). The other approach, known as Diffusion Classifiers, repurposes diffusion models for zero-shot classification using their original training objective Li et al. (2023); Krojer et al. (2023); Clark and Jaini (2023); He et al. (2024); Jaini et al. (2023). Notably, the latter has outperformed CLIP on compositional benchmarks like CLEVR Johnson et al. (2017) and Winoground Thrush et al. (2022), which require reasoning over multiple objects and attributes. However, existing studies are limited in scope, often relying on a small number of benchmarks and lacking a systematic analysis.
@@
Hypothesis 3: The domain gap can be mitigated by timestep weighting. Lastly, we investigate how diffusion timesteps impact discriminative performance by optimizing timestep weights for downstream tasks. Prior work has explored how diffusion models generate images through a structured progression across timesteps Li et al. ; Wang and Vastola (2023); Kim et al. (2024); Zheng et al. (2024a); Wang et al. (2025). However, in diffusion classifiers, either uniform timesteps or fixed timestep weightings are typically used across all datasets and models—an area that remains underexplored. In contrast, we find that SD3-m is particularly sensitive to timestep selection. Our results show that even low-shot timestep tuning (using just 5% of the target dataset can significantly mitigate the performance drop of SD3-m on real-world compositional benchmarks). We hypothesize that SD3-m’s timestep sensitivity is closely linked to its susceptibility to the domain gap. To further examine this relationship, we incorporate CLIP-based image encoders to quantify visual similarity between domains, and analyze how it correlates with optimal timestep weighting. We find that timestep weighting is especially effective in scenarios with large domain gaps.
@@ 2 Related work
Diffusion models. Generative models have demonstrated impressive performance in producing realistic images Ramesh et al. (2022); Yang et al. (2024); Chen et al. (2023); Yu et al. (2024), videos Blattmann et al. (2023); Zheng et al. (2024b); Jeong et al. (2023), and audio Evans et al. (2024); Liu et al. (2023a); Jeong et al. (2025). In particular, text-to-image diffusion models Ho et al. (2020) iteratively refine images conditioning on text prompts by adding and removing noise, achieving remarkable quality. A widely used open-source example is Stable Diffusion (SD) Rombach et al. (2022). Earlier SD versions Rombach et al. (2022) are based on a UNet Ronneberger et al. (2015) backbone, incorporating ResNet blocks He et al. (2016) and attention mechanisms Vaswani et al. (2017). With the rise of transformer-based designs Vaswani et al. (2017), the latest Stable Diffusion 3 series Esser et al. (2024) adopts this architecture, further boosting performance. It also introduces new noise sampling strategies for training Rectified Flow models Liu et al. (2023b); Albergo and Vanden-Eijnden (2023); Lipman et al. (2023). Our analysis focuses on Stable Diffusion versions 1.5, 2.0, and 3-m, examining their architectural evolution and performance across scales.
@@ 3.1 Preliminaries: diffusion classifiers
where $w_{t}$ are timestep weights, $\epsilon_{\Theta}$ is a neural network that predicts the noise $\boldsymbol{\epsilon}$ added to the latent $\mathbf{z}$ at timestep $t\in T$, and $\mathbf{c}$ is a conditioning text prompt. Unless stated otherwise we draw $t\sim\text{Uniform}([0,1])$. This loss is related to the ELBO of the conditional likelihood $p(\mathbf{z}|y)$, where $y$ is the class label, which allows us to use diffusion models for classification, as shown in Li et al. (2023); Krojer et al. (2023); Clark and Jaini (2023):
@@ 3.2 SD3-m as a classifier
SD3-m is a rectified flow model Liu et al. (2023b); Albergo and Vanden-Eijnden (2023) trained with a conditional flow matching (CFM) loss Lipman et al. (2023), which differs from the standard diffusion objectives used in SD1.5 and SD2.0. As a result, we cannot directly apply the same classifier objective used in earlier versions. By reparameterizing the CFM objective as a noise‐prediction loss (see, e.g., Esser et al. (2024)), we can obtain
@@
Using this formulation, we can use SD3 as classifiers, despite its different underlying architecture. The only difference lies in the weighting function $w_{t}$, which for SD3 follows a logit-normal distribution rather than the uniform weighting used in SD1.5/2.0. However, we empirically find that uniform weighting performs better for classification. Details are given in Section A.
@@ 4 Self-Bench: a diagnostic benchmark
As shown in Figure 2 (left), existing benchmarks are diverse in terms of domains. However, we empirically observe that the generation results of diffusion models are not as diverse, having a preferred “native” style. For example, in SD3-m, objects are usually well-centered/focused and have a glossy aesthetic touch; images are in high resolution (see Figure 2, right). Additional examples in Figure D.3 (Supplemental) further support that SD3-m consistently produces images with a similar style.
@@
Figure 3: Diagnosing with Self-Bench. (i) Using Geneval’s prompts from six categories, generate images. (ii) For each generated image, create discriminative tasks within its type from the prompts used in the generation process. (iii) Given the generated images (filtered by humans) and the discriminative tasks, benchmark the diffusion classifier.
@@ 5.1 Experimental setting
Stable Diffusion baselines. For the evaluation of diffusion classifiers, we use three versions of Stable Diffusion models: SD1.5, SD2.0 Rombach et al. (2022), and SD3-m Esser et al. (2024); each new version increases the number of parameters. We selected the baselines based on specific criteria, which are detailed in Section C.1 of the Supplemental. Although we also considered distilled variants of diffusion models (e.g., SDXL-Turbo Sauer et al. (2024)), we excluded them from our main evaluation due to their architectural and generative differences. Additional analysis is provided in Section E.2 of the Supplemental. For SD1.5 and SD2.0, we use the Euler Discrete scheduler Karras et al. (2022) and uniformly sample the timesteps. For SD3-m, we use the Flow Matching EulerDiscrete scheduler Esser et al. (2024) for flow matching diffusions, which was designed specifically for the SD3 series. We sample 30 timesteps from each model uniformly, following He et al. (2024). (The effect of different numbers of timesteps is discussed in Table B.2 in the Supplemental.)
CLIP baselines. We use vanilla CLIP models as key baselines for comparison on discriminative tasks. We use five different versions of CLIP models: RN50x64, ViT-B/32, ViT/L14, ViT/H14, and ViT/g14. We follow OpenAI’s implementation for the first three: RN50x64, ViT/B32, and ViT/L14 Radford et al. (2021). On the other hand, we follow OpenCLIP’s implementation for ViT-H/14 and ViT-g/14 Ilharco et al. (2021). Where appropriate, we also report SigLIP and SigLIP2 for completeness, but their behavior closely mirrors that of CLIP.
Metrics. The benchmarks vary in structure. Some present paired image-text inputs (i.e., two images and two prompts), while others use a single image with multiple candidate prompts. Across all setups, we primarily evaluate using image-to-text retrieval accuracy, which measures whether the model assigns the highest score to the correct prompt given an image. For paired settings, we compute retrieval accuracy based on whether the matching image-prompt pairs are correctly ranked relative to distractors. Further details on task-specific evaluation protocols are provided in Section C.2.
@@ 5.2 Scaling evaluation to ten benchmarks
Benchmarks. Our analysis incorporates ten benchmarks: Vismin Awal et al. (2024), EQBench Wang et al. (2023), MMVP Tong et al. (2024), CLEVR Johnson et al. (2017), Whatsup Kamath et al. (2023), Spec Peng et al. (2024), ARO Yuksekgonul et al. (2023), Sugarcrepe Hsieh et al. (2023), COLA Ray et al. (2023), and Winoground Thrush et al. (2022). Each benchmark contains different tasks. For example, Vismin includes tasks related to Object, Attribute, Position, and Counting, whereas COLA is focused on Attribute tasks. The left block in Figure 2 presents examples of these benchmarks. Further details on how the tasks are classified within our compositional task categories are provided in Table C.1 in the Supplemental. In total, we analyze 33 tasks in our main study.
@@
Results. Figure 4 shows the average performance of diffusion classifiers on the compositional benchmarks (see Figure E.1 in the Supplemental for the complete results). [^2] Tables E.1E.2E.3E.4E.5E.6E.7E.8E.9 and E.10 report the quantitative results for all benchmarks used in our evaluation. For the Position task, SD3-m performs the best compared to other diffusion models and CLIP models. However, in other tasks, CLIP models usually show better performance than diffusion classifiers, contrary to the findings of previous works Li et al. (2023); Clark and Jaini (2023). Similar results hold for SigLIP (see E.7 in Appendix). Surprisingly, among diffusion classifiers, SD3-m is not the best; often SD1.5 or SD2.0 models show better results. These results motivate us to deepen our analysis of when and why diffusion classifiers may underperform.
@@ 5.3 Studying domain effects via Self-Bench
In-domain evaluation. First, we examine how diffusion classifiers behave in-domain, which may serve as an upper-bound performance estimate, as domain effects are isolated. As shown in Figure 5, they perform better on Correct (blue bars) than on Full (bright yellow bars) samples. This indicates that the classifiers do not simply choose prompts used for generation; rather, they indeed distinguish between the correct and incorrect prompts.
@@
Figure 6 illustrates the average accuracy drop when moving from in-domain to cross-domain evaluation on Self-Bench. We observe accuracy degradation across all tasks, with SD3-m showing the most drop. While part of the performance gap may be due to task-specific weaknesses in each model, we assume the domain gaps, such as different data distribution, also play a key role in this degradation. Full results can be found in Figure E.7 in the Appendix.
@@ 5.4 Timestep weighting effects
Previous works have shown that diffusion models generate images from coarse to fine details over timesteps Li et al. ; Wang and Vastola (2023). However, in classification settings, it is still unclear how different noise levels (i.e., timesteps) affect performance across tasks (e.g., object recognition vs. attribute binding) or different domains (e.g., image style). Diffusion classifiers typically use uniform timestep weighting Li et al. (2023); Krojer et al. (2023) or a fixed timestep weighting scheme (e.g., $w_{t}=\exp(-7t)$Clark and Jaini (2023); Jaini et al. (2023) across all models (e.g., Imagen Saharia et al. (2022) and SD). In generation, however, recent works have shown that non-uniform timestep sampling can substantially affect sample quality and training dynamics Kim et al. (2024); Zheng et al. (2024a); Wang et al. (2025).
We hypothesize that neither strategy is universally optimal (see ablation studies of uniform weighting and fixed timestep weighting in Sec. B.3 in the Supplemental) and that timestep weighting should be adapted to the model and task differently. Here, we investigate how different timesteps contribute to classification decisions. (Figure 10 in the Supplemental provides an intuitive explanation of why timesteps matter from a generative perspective.)
@@
To study this further, we conduct an experiment to approximate the effect of domain differences. We generate images using the original prompts from real-world compositional benchmarks. This yields two image sets for each task: (i) the original real-world dataset, and (ii) a synthetic variant generated using the same prompts. Both sets target the same task but differ in visual domain. Using a CLIP image encoder (ViT-B/32) Radford et al. (2021), we aim to capture the domain gap [^3] Indeed CLIP may capture both stylistic as well as semantic shifts, broadly referred to as ”domain.” between the two, by computing the L2 distance between average embeddings with randomly sampled 50 images. The datapoints in Figure 9 are based on the same real-world benchmarks used in Figure 8. As shown, for SD3, larger CLIP embedding distances (i.e., greater domain gap) are associated with greater reliance on timestep weighting. This suggests a positive correlation between domain gap and the effectiveness of timestep weighting for SD3. In contrast, SD1 and SD2 do not exhibit such a trend. We hypothesize that, as shown in Figure 7, most timesteps in SD2 (also in SD1) are already effective, resulting in limited benefit from reweighting and thus obscuring any correlation with the domain gap. Details are in Section E.6 of the Appendix.
@@
To build intuition for how timestep selection impacts classification, we visualise generations from the SD3-m model in Figure 10, starting from a Self-Bench image generated by SD2.0 (“a parking meter and a teddy bear”). For each timestep $t$, we corrupt the original image with Gaussian noise corresponding to $t$, and then generate an image by running the diffusion model for 20 denoising steps from $t$ down to 0, conditioning on either a correct or incorrect prompt. We set the classifier-free guidance coefficient to 0.0 (i.e., using only the conditional prompt) to match the discriminative setting used in classification with diffusion.